OpenStudy (anonymous):

Find the inverse of the function. f(x)=x/7−−−√3−9

OpenStudy (anonymous):

f(x)=√x/7 -9

OpenStudy (anonymous):

@bridgetx516x @cool1939 @divagirl421 @feezul_93 @Frustratedashell

OpenStudy (anonymous):

i believe u should make ur question {equation} more clear.

OpenStudy (anonymous):

\[f(x)= \sqrt{x/7} -9\]

OpenStudy (anonymous):

much better :D

OpenStudy (anonymous):

its a cubed root sorry

OpenStudy (anonymous):

if its f(x)=(x/7)^(1/3) -9, then its inverse is f(x)=7(x+9)^3